40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Dec 10 15:05:32 2014, Jenne, Update, Electronics, Xend QPD whitening board modified already IMG_1776.JPG
    Reply  Thu Dec 11 16:42:12 2014, Jenne, Update, Electronics, Xend QPD whitening board plan QPDwhiteningModification_11Dec2014.pdf
       Reply  Fri Dec 12 19:54:21 2014, Jenne, Update, Electronics, Xend QPD whitening board modified QPDwhiteningModification_XtransCompleted_12Dec2014.pdfIMG_1779.JPGQPD_Xtrans_Ytrans_12Dec2014.pdfQPD_Xtrans_Ytrans_DarkVsLock_12Dec2014.pdf
          Reply  Sat Dec 13 00:35:11 2014, rana, Update, Electronics, Xend QPD whitening board modified 
          Reply  Mon Dec 15 22:30:50 2014, Jenne, Update, Electronics, Yend QPD modified 
             Reply  Mon Dec 15 22:45:59 2014, Jenne, Update, Electronics, Yend QPD modified QPD_Ytrans_oscillating_15Dec2014.pdfQPD_IOPchannels_Ytrans_oscillating_15Dec2014.pdf
                Reply  Fri Feb 13 14:27:30 2015, ericq, Update, Electronics, Second QPD Whitening Switch enabled 
                   Reply  Fri Feb 13 14:59:13 2015, ericq, Update, Electronics, Second QPD Whitening Switch enabled 
                      Reply  Fri Feb 13 18:56:44 2015, rana, Update, Electronics, Second QPD Whitening Switch enabled 
                         Reply  Thu Feb 19 04:09:21 2015, ericq, Update, Electronics, Second QPD Whitening Switch enabled 
                   Reply  Fri Feb 13 17:07:51 2015, Jenne, Update, Electronics, Second QPD Whitening Switch enabled 
Message ID: 11049     Entry time: Thu Feb 19 04:09:21 2015     In reply to: 11025
Author: ericq 
Type: Update 
Category: Electronics 
Subject: Second QPD Whitening Switch enabled 

X end QPD has recieved 0.2+0.4 absorptive ND filters. Y end QPD got one at 0.6. This appears to have mitigated the saturations for now; the unwhitened signals no longer go negative. The digital gains have been reset. 

 

ELOG V3.1.3-